OpenStudy (anonymous):

tan(2Θ)+tan(Θ)= 0 SOLVE ON THE INTERVAL: 0≤Θ<2π

OpenStudy (anonymous):

@mustafa2014 help pls?

OpenStudy (anonymous):

\[\tan(2Θ)+\tan(Θ)= 0\] \[\frac{ 2\tan(Θ) }{ 1-\tan^2(Θ) }+\tan(Θ)=0\] \[\frac{ 2\tan(Θ) }{ 1-\tan^2(Θ) }+\frac{ tan(Θ)(1-\tan^2(Θ) }{ 1-\tan^2(Θ) }=0\] \[\frac{ 2\tan(Θ)+tan(Θ)-tan^3(Θ) }{ 1-\tan^2(Θ) }=0\] \[3\tan(Θ)-tan^3(Θ)=0\]

OpenStudy (anonymous):

wait so after 3tanΘ and all that do u factor

OpenStudy (anonymous):

\[\tan(Θ)(3-tan^2(Θ)) =0\]

OpenStudy (anonymous):

and then u solve for each individually right?

OpenStudy (anonymous):

like make both= 0 and solve?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

is that the right way though? for sure?

OpenStudy (anonymous):

thank you again.

OpenStudy (anonymous):

wut

OpenStudy (anonymous):

wait no thats not how youd solve. thats only if theyre muliplying each other

OpenStudy (anonymous):

@Mertsj does this seem right ? it seems wrong somehow?

OpenStudy (mertsj):

Is the original problem tan(2x) or tan^2x

OpenStudy (anonymous):

its tan 2theta not squared

OpenStudy (mertsj):

\[3\tan \theta -\tan ^3\theta =0\] \[\tan \theta (3-\tan ^2\theta)=0\] \[\tan \theta=0; or\] \[3-\tan ^2\theta=0\]

OpenStudy (mertsj):

Solve each equation

OpenStudy (anonymous):

where r u getting the 3 from??

OpenStudy (anonymous):

the original equaion is tan(2Θ)+tan(Θ)= 0

OpenStudy (mertsj):

Did you not follow the work that mustafa posted above?

OpenStudy (mertsj):

He substituted for tan 2theta and simplified for you. He even did the factoring for you. All you had to do was set each factor equal to 0 and solve So I set each factor equal to 0 Now all you have to do is solve.

OpenStudy (mertsj):

If you do not trust his work (and I did not check it) all you have to do is replace tan 2theta with 2tan(theta)/1-tan^2(theta) and simplify for yourself.

OpenStudy (anonymous):

yes but someone else helped me w/ same problem and the final answer they got was Θ= sqrt 3 and then you got the solutions.

OpenStudy (mertsj):

theta is an angle. If they got tan theta = sqrt 3, that would make sense because then theta would be 60 degree. But theta = sqrt 3 is suspect.

OpenStudy (anonymous):

o=sorry it was tanΘ=sqrt 3 so is that the same thing as what mustafa said?

OpenStudy (anonymous):

wait it is right? they both said the same thing then?